### v2

# Pseudo Twirling Mitigation of Coherent Errors in non-Clifford Gates

Abstract
The conventional circuit paradigm, utilizing a limited number of gates to construct arbitrary quantum circuits, is hindered by significant noise overhead. For instance, the standard gate paradigm employs two CNOT gates for the partial CPhase rotation in the quantum Fourier transform, even when the rotation angle is very small. In contrast, some quantum computer platforms can directly implement such operations using their native interaction, resulting in considerably shorter and less noisy implementations for small rotation angles. Unfortunately, coherent errors stemming from qubit crosstalk and calibration imperfections render these implementations impractical. In Clifford gates such as the CNOT, these errors can be addressed through Pauli twirling (also known as randomized compiling). However, this technique is not applicable to the non-Clifford native implementations described above. The present work introduces, analyzes, and experimentally demonstrates a technique called ‘Pseudo Twirling’ to address coherent errors in general gates and circuits. Additionally, we experimentally showcase that integrating pseudo twirling with a quantum error mitigation method called ‘Adaptive KIK’ enables the simultaneous mitigation of both noise and coherent errors in non-Clifford gates. Due to its unique features pseudo twirling could become a valuable asset in enhancing the capabilities of both present and future NISQ devices.
I Introduction
In recent years, quantum error mitigation protocols (QEM) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] have significantly boosted the performance of quantum computers [11, 12, 5, 13, 14, 15, 16, 17, 18, 19]. These methods are primarily designed for managing incoherent errors (noise) arising from interactions with the environment. However, in addition to incoherent errors, there are also coherent errors that cannot be addressed using the same methods in a scalable manner. Coherent errors pose a significant concern in the development of quantum computers in the foreseeable future. They are often attributed to calibration errors, drift in calibration parameters, or coherent crosstalk interactions between qubits. While hardware continues to advance, the demand for minimizing coherent errors becomes increasingly crucial as circuits grow in size and complexity.
Currently, the most effective approach to address these errors is a method known as Randomized Compiling (RC) [20, 15, 21, 22, 23], also referred to as Pauli twirling. However, this technique is limited to Clifford gates. For Clifford gates such as the CNOT, iSWAP, and others, RC involves executing an ensemble of circuits that are equivalent in the absence of coherent errors, but each circuit alters the coherent error in a distinct way. After running these circuits and performing post-processing averaging, the coherent error is mitigated, but the resulting transformation is no longer a unitary. In other words, the coherent error is converted into an effective incoherent error. This conversion is beneficial as incoherent errors can be addressed with QEM protocols. As demonstrated experimentally in [24] the integration of RC with a QEM method called ‘Adaptive KIK’, was imperative for removing coherent errors and obtain accurate results.
Another interesting feature of RC is that it converts a general incoherent error channel into a Pauli channel that has a much simpler structure [20]. Due to the reduced degrees of freedom in the Pauli channel compared to general noise, it becomes feasible to employ additional sparsity techniques and efficiently characterize the noise channel of multi-qubit Clifford gates [18, 25, 21]. This key feature facilitates two important QEM techniques, PEC [4] and PEA [11]. PEA presently holds the record of the largest circuit that QEM has been applied to – 127 qubit Ising lattice time-evolution experiment [11]. The largest depth probed in this experiment was sixty cnots.
Unfortunately, RC is restricted to Clifford gates, rendering PEC and PEA inapplicable to non-Clifford multi-qubit gates. A recent work [26] extends these methods to non-Clifford gates. In contrast to the Clifford case, the sampling overhead per gate can be substantial regardless of how small the coherent error is. In [27] a partial RC was applied by using only the Paulis that commutes with the non-Clifford gate. This partial twirling can only address coherent errors that happen to anti-commute with the selected Pauli subset. Thus, it is not possible to rely on this method if the coherent error is of unknown type, or if it known to commute with the selected Pauli subset.
Circuits consisting solely of Clifford gates can be efficiently simulated, making it pointless to build a powerful quantum computer just for their execution. Although multi-qubit non-Clifford gates are not part of the conventional universal gate set paradigm, they can prove crucial in specific applications, particularly in the era of noisy quantum computers where circuit depth is a limiting factor.
As a first example, we bring the time-evolution of the transverse Ising lattice model [28]. The circuits implementing this evolution require $R_{zz}(\theta)$ gates that can be directly implemented very rapidly, leading to significantly lower noise levels. For example, in [27] the IBM cross-resonance interaction [29, 30, 31], was facilitated to directly implement $R_{zz}(\theta)$ using two additional single qubit gates as depicted in Fig. 1a. In contrast, in the cNOT gate paradigm, the length of the ZZ rotation gate is determined by the length of two CNOT gates, even when the rotation angle is minute (see 1b). The direct implementation in [27] has enabled the execution of Ising time evolution that are four times longer compared to the double-CNOT implementation (Fig. 1b). In trapped ions, a similar statement holds for the Mølmer–Sørensen gate [32]. Although in this case the direct non-Clifford gate may not be shorter than a single CNOT gate, it is at least twice shorter than the double-CNOT implementation.
The second example pertains to the $n$-qubit quantum Fourier transform [33] that comprises of single-qubit gates and two-qubit gates that execute CPhase($\theta$) rotations with an angle of $\theta=\pi/2^{n}$ (and multiples of it). As $n$ increases, the gate performs a transformation that is very close to the identity transformation. As such, a substantial advantage in terms of noise is expected when using the direct implementation (the CPhase($\theta$) analogue of Fig. 1a) instead of the double-CNOT implementation. Hence, a method for mitigating coherent errors in non-Clifford gates would benefit the implementation quantum Fourier transforms and other circuits as well.
Figure 1: Non-Clifford vs. Clifford implementation of quantum gates. (a) The useful two-qubit gate $R_{zz}(\theta)=e^{-i\frac{1}{2}\theta Z_{1}Z_{2}}$ can be implemented compactly in the IBM platform using a non-Clifford gate $R_{zx}(\theta)=e^{-i\frac{1}{2}\theta Z_{1}X_{2}}$, generated by the cross-resonance interaction. Conversely, the current Clifford-based implementation (b) is substantially longer, with each box representing a CNOT gate. Consequently, this implementation exhibits considerably higher levels of noise. However, currently, there is no technique for addressing coherent errors in non-Clifford gates, providing an advantage to the more noisy implementation depicted in (b). In this work we present a protocol for mitigating coherent errors in compact, low-noise non-Clifford gates, as shown in (a).
In this work we present a new twirling scheme called ‘Pseudo Twirling’ (PST). This approach offers three noteworthy characteristics: 1) PST is equally applicable to both Clifford and non-Clifford gates. 2) In contrast to RC (Pauli twirling), PST distinguishes between controlled over(under)-rotation errors and all other controlled or uncontrolled coherent errors. Over-rotations typically arise from imperfect initial calibration or a calibration drift. Others errors stem from various mechanism ranging from crosstalk between qubits to parasitic higher-order interactions created the driving fields. PST transforms all coherent error excluding controlled rotation errors - a fact that can be exploited for high accuracy calibration. 3) PST does not produce a Pauli error channel; instead, it simplifies the noise in a different manner. In this paper we present and study these three features. The integration of PST with QEM, as demonstrated in the present work, can have an immediate impact on the capabilities of quantum computers.
With the restriction of applying twirling exclusively to Clifford gates removed, PST can be employed within individual components of a gate (intra-gate PST). As we show later, this facilitates more effective mitigation of coherent errors since coherent errors are addressed before they have a chance to accumulate. Additionally, PST’s flexibility allows for its application across an entire circuit, where the twirling occurs at the edges. Although this ‘Edges PST’ is less effective than intra-gate PST, it can capture coherent errors occurring on a scale exceeding that of multi-qubit gates.
Notations and quantum dynamics in Liouville space
For convenience we shall use Liouville space representation where the density matrix is flattened into a column vector $\left|\rho\right\rangle$ and the evolution is described by superoperators operating on this vector from the left $\left|\rho(t)\right\rangle=\mathcal{K}(t)\left|\rho(0)\right\rangle$. If the dynamics is unitary the evolution of the density matrix $\left|\rho(t)\right\rangle$ is determined by Schrödinger-like equation where
$\frac{d}{dt}\left|\rho(t)\right\rangle$ $=-i\mathcal{H}(t)\left|\rho(t)\right\rangle,$ (1)
$\mathcal{H}(t)$ $=H(t)\otimes I-I\otimes H(t)^{\intercal}.$ (2)
[…]
$\mathcal{U}(t)=U(t)\otimes U(t)^{*},$ (3)
where $U(t)$ is the evolution operator in Hilbert space. Pauli matrices in Liouville space can either appear as Hamiltonians or as unitaries, but unlike in Hilbert space, these two forms differ. Let us denote by $P_{\alpha}$ the tensor product of single-qubit Pauli matrices $\sigma_{i}$ such that $P_{\alpha}\in\{\sigma_{k}\otimes\sigma_{l}\otimes\sigma_{m}...\}_{k,l,m..\in\{0,x,y,z\}}$. A Pauli Hamiltonian in Liouville space is given by $\mathcal{H}_{\alpha}=P_{\alpha}\otimes I-I\otimes P_{\alpha}^{\intercal}$, while according to Eq. (3), a Pauli evolution operator has the form $\mathcal{P}_{\alpha}=P_{\alpha}\otimes P_{\alpha}^{*}$ in Liouville space. These two are related via $\mathcal{P}_{\alpha}=exp(-i\frac{\pi}{2}\mathcal{H}_{\alpha})$. See Appendix I for more details.
II Description of the PST methodology
II.1 Regular Randomized Compiling
For clarity, we start by describing the regular randomized compiling/Pauli twirling scheme, which is applicable only for Clifford gates. Let $\mathcal{U}_{\text{cliff}}$ stand for the evolution operator of an ideal Clifford gate, and let $\mathcal{K}_{\text{cliff}}$ stand for the evolution operator of the noisy realization of $\mathcal{U}_{\text{cliff}}$. $\mathcal{K}_{\text{cliff}}$ may encompass both noise and coherent errors, but, for the moment, we will concentrate on the noise-free scenario where only coherent errors take place.
As shown in Fig. 2a, in RC, $\mathcal{K}_{\text{cliff}}$ is replaced by an averaged operator, with each element “twirled” by an n-qubit Pauli operators $\mathcal{P}_{\alpha}$ and $\mathcal{P}^{\prime}_{\alpha}$
$\mathcal{K}_{\text{cliff}}^{RC}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}^{\prime}\mathcal{K}_{\text{cliff}}\mathcal{P}_{\alpha}.$ (4)
$\mathcal{P}^{\prime}_{\alpha}$ is chosen in such a way that it yields $\mathcal{K}_{\text{cliff}}^{RC}=\mathcal{U}_{\text{cliff}}$ in the absence of coherent and incoherent errors. This condition of invariance can be expressed as $\mathcal{P}_{\alpha}^{\prime}\mathcal{U}_{\text{cliff}}\mathcal{P}_{\alpha}=\mathcal{U}_{\text{cliff}}$, leading to the determination of $\mathcal{P}_{\alpha}^{\prime}$ as $\mathcal{P}_{\alpha}^{\prime}=\mathcal{U}_{\text{cliff}}\mathcal{P}_{\alpha}\mathcal{U}_{\text{cliff}}^{\dagger}$. Therefore, $\mathcal{P}_{\alpha}^{\prime}$ is determined by the selected $\mathcal{P}_{\alpha}$. Up to his point it appears as if we have not used the fact that $\mathcal{U}_{\text{cliff}}$ is a Clifford gate. However, since Clifford gates normalize the Pauli operators, it is guaranteed that if $\mathcal{P}_{\alpha}$ is a Pauli operator, then $\mathcal{P}_{\alpha}^{\prime}$ is also a Pauli operator. Conversely, if the ideal gate is non-Clifford, $\mathcal{P}_{\alpha}^{\prime}$ may involve multi-qubit interactions, introducing noise and coherent errors comparable to those RC aims to correct. These additional errors could potentially degrade the performance of the gate with respect to the non-twirled gate. Thus, a significant advantage of Pauli Twirling lies in the experimental simplicity of the twirling operations, which involve very low levels of coherent and incoherent errors.
Notably, the total number of twirlings for $n$-qubit gate, $2^{2n}$, can swiftly grow to a very substantial value. This becomes particularly evident when considering sequences of twirled gates. Fortunately, there is no need to sum over all possible Pauli operators. It suffices to sample from the set $\{\mathcal{P}_{\alpha}\}{}_{\alpha=1}^{2^{2n}}$ to obtain the average value. Nevertheless, for the sake of simplicity of the derivations we will maintain the use of $2^{2n}$.
To gain a better grasp of how RC operates, let us express $\mathcal{K}_{\text{cliff}}$ as $\mathcal{K}_{\text{cliff}}=\mathcal{U}_{\text{cliff}}\mathcal{N}$, where $\mathcal{N}$ represents the error channel that encompasses all errors. This decomposition always exists when the dynamics of the gate constitute a completely positive map. Substituting it into the twirling formula, we obtain:
$K_{\text{cliff}}^{RC}$ $=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}^{\prime}\mathcal{U}_{\text{cliff}}\mathcal{N}\mathcal{P}_{\alpha}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}^{\prime}\mathcal{U}_{\text{cliff}}\mathcal{P}_{\alpha}\mathcal{P}_{\alpha}\mathcal{N}\mathcal{P}_{\alpha}$
$=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{U}_{\text{cliff}}\mathcal{P}_{\alpha}\mathcal{N}\mathcal{P}_{\alpha}=\mathcal{U}_{\text{cliff}}\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{N}\mathcal{P}_{\alpha}.$ (5)
Hence, twirling a noisy Clifford gate with $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\alpha}^{\prime}$ is equivalent to twirling its noise channel $\mathcal{N}$ with the same $\mathcal{P}_{\alpha}$ on both sides ($\mathcal{P}_{\alpha}^{\prime}=\mathcal{P}_{\alpha}$). Furthermore, we can utilize the fact that Pauli operators form a complete basis, allowing to express any $\mathcal{N}$ as $\mathcal{N}=\sum_{\alpha,\beta}\mathcal{N}_{\alpha,\beta}P_{\alpha}\otimes P_{\beta}^{*}$ It can be shown that, after Pauli twirling, $\mathcal{N}$ transforms into
$\mathcal{N}^{RC}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{N}\mathcal{P}_{\alpha}=\sum_{\alpha}\mathcal{N}_{\alpha,\alpha}\mathcal{P}_{\alpha}.$ (6)
In this transformation, only the diagonal elements remain. This resulting channel is referred to as a Pauli error channel, and it corresponds to randomly applying a Pauli unitary $\mathcal{P}_{\alpha}$ with a probability of $\mathcal{N}_{\alpha,\alpha}$. Thus, regardless of the original structure of $\mathcal{N}$, Pauli twirling turns it into a Pauli channel. As we show later, this is not the case in PST.
Figure 2: (a) Randomized compiling (RC, AKA Pauli twirling) replaces the original Clifford (e.g. a CNOT) with a sum of Pauli-twirled Clifford gates. RC can only mitigate coherent errors in Clifford gates. (b) Our proposed Pseudo Twirling (PST) scheme can mitigate uncontrolled coherent errors, such as crosstalk, in a general gate. This is achieved by modifying the signs of some of the pulses in the driving field $H_{d}(t)$ that generates the gate according to the rule $H_{d}(t)\to P_{\alpha}H_{d}(t)P_{\alpha}$. (c) Because PST is not restricted to Clifford gates, it can be implemented at various levels of a circuit: at the entire circuit level (orange lines), gate-level (green lines), and intra-gate level (purple lines).
II.2 Introducing the Pseudo Twirling Protocol
To introduce the concept of pseudo twirling we start with the following example. Consider a simple non-Clifford gate, such as the two-qubit gate $\mathcal{U}_{zz}(\theta)=e^{-i\theta\mathcal{H}_{zz}}$, with $\theta\neq k\pi/2$. For simplicity, we use the notation $\mathcal{H}_{zz}=P_{zz}\otimes I-I\otimes P_{zz}^{\intercal}$ , $P_{zz}=\sigma_{z}\otimes\sigma_{z}$. In analogy to Pauli twirling we aim to create an operation that keeps the functionality of the ideal gate. If we apply $\mathcal{P}_{xx}=P_{xx}\otimes P_{xx}^{*}$, we get: $\mathcal{P}_{xx}e^{-i\theta\mathcal{H}_{zz}}\mathcal{P}_{xx}=e^{-i\theta\mathcal{H}_{zz}}$ since $\mathcal{P}_{xx}$ and $\mathcal{H}_{zz}$ commute. However, $\mathcal{P}_{xz}$, for example, anti-commutes with $\mathcal{H}_{zz}$, and therefore $\mathcal{P}_{xz}e^{-i\theta\mathcal{H}_{zz}}\mathcal{P}_{xz}=e^{+\imath\theta\mathcal{H}_{zz}}=\mathcal{U}_{zz}(-\theta)$. Our PST protocol is based on the fact that if the Pauli operators anti-commute with the Hamiltonian of $U$ (or equivalently $\mathcal{U}$ in Liouville space), we can simply change the sign of the angle, i.e., the sign of the driving fields, to achieve the desired transformation in the end:
$\mathcal{P}_{xz}\mathcal{U}_{zz}(-\theta)\mathcal{P}_{xz}=\mathcal{U}_{zz}(\theta).$ (7)
More generally when the driving Hamiltonian is not a single Pauli, we can write $\mathcal{U}=e^{-i\sum h_{\beta}\mathcal{H}_{\beta}}$ and then the PST twirling takes the form
$\mathcal{K}_{PST}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}e^{-i\mathcal{P}_{\alpha}(\sum_{\beta}h_{\beta}\mathcal{H}_{\beta})\mathcal{P}_{\alpha}}\mathcal{P}_{\alpha}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}e^{-i\sum_{\beta}sgn(\alpha,\beta)h_{\beta}\mathcal{H}_{\beta}}\mathcal{P}_{\alpha},$ (8)
where $sgn(\alpha,\beta)=tr(P_{\alpha}P_{\beta}P_{\alpha}P_{\beta})/2^{n}$ equals $+1$ if $P_{\alpha}$ and $P_{\beta}$ commute, and $-1$ if they anti-commute. Note that in the definition of $sgn(\alpha,\beta)$ we used Pauli matrices in Hilbert space. Mathematically, expression (8) is straightforward, but it should be considered with execution protocol in mind. The right $\mathcal{P}_{\alpha}$ is executed at the beginning of the gate. Then a modified unitary $\mathcal{U}_{\alpha}=e^{-i\sum_{\beta}sgn(\alpha,\beta)h_{\beta}\mathcal{H}_{\beta}}$ is executed next by changing the sign of some of the control signals. Finally, another $\mathcal{P}_{\alpha}$ is executed at the end of the gate. Importantly, the Hamiltonian of the modified unitary has the same elements as the original Hamiltonian, with only the sign of some terms changing. Thus, under the assumption that the signs of the driving Hamiltonian terms are controllable, the PST implementation is no more challenging than implementing the original unitary.
The main operational differences between RC and pseudo twirling are:
  • In RC, the Pauli gates at the beginning and at the end are different. In pseudo twirling, they are always the same.
  • In RC, the operation between the Pauli gates (the original gate) is always the same. Conversely, in PST, the operation between the Pauli gates depends on the selected Pauli.
  • In both methods, the twirling gates are Pauli gates, but in PST, the Pauli gates are also employed to determine the signs of the control signals that drive the gate.
RC and PST are illustrated in Fig. 2a and Fig. 2b respectively. We note that the term ‘pseudo twirling’ reflects not merely an operational difference from standard twirling but also its distinct impact on errors, as discussed in the following two sections. Its effects differs from Pauli twirling in two significant ways: i) it cannot eliminate over/under rotations errors, and ii) in general, it does not transform incoherent errors into a Pauli channel. Nevertheless, we show and demonstrate that these differences do not diminish the usefulness of pseudo twirling.
II.3 PST Impact on Coherent Errors
In this section we study the impact of PST on coherent errors. Through the Magnus expansion [34] in the interaction picture, we find that the application of PST in a circuit with a coherent error induced by a Hamiltonian term $\mathcal{H}_{coh}(t)$ leads to the following evolution operator
[…]
where $\mathcal{U}(t)$ is the ideal evolution operator of the non-twirled gate and $\mathcal{U}_{T}=\mathcal{U}(T)$. The second and higher order terms in the Magnus expansion have been neglected. $\mathcal{H}_{coh}$ is in Liouville space which is related to the Hilbert space Hamiltonian $H_{coh}$ through Eq. (2). The first order Taylor expansion in $\mathcal{H}_{coh}$ is
[…]
$\mathcal{U}_{T}+\mathcal{U}_{T}\frac{-i}{2^{2n}}\intop_{0}^{T}\mathcal{U}(t)^{\dagger}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{H}_{coh}\mathcal{P}_{\alpha}\mathcal{U}(t)dt+O(\mathcal{H}_{coh}^{2})$ $=\mathcal{U}_{T}+O(\mathcal{H}_{coh}^{2}),$ (10)
where the last step follows from the fact that $\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{H}_{coh}\mathcal{P}_{\alpha}$ is a Pauli twirling of $\mathcal{H}_{coh}$. As such, the twirling eliminates the leading term in the coherent error since $\sum_{\alpha=1}^{2^{2n}}P_{\alpha}H_{coh}P_{\alpha}=tr(H_{coh})I$ and $\sum_{\alpha=1}^{2^{2n}}P_{\alpha}^{*}H_{coh}^{\intercal}P_{\alpha}^{*}=tr(H_{coh})^{*}I=tr(H_{coh})I$ so that the two terms in $\mathcal{H}_{coh}$ cancel each other. This explains why PST suppresses coherent errors.
Next, we study the effect of PST on the second order Taylor approximation in $\mathcal{H}_{coh}$. This term is important in light of the fact the first order is eliminated by PST. Furthermore, this analysis will unravel the structure of the resulting noise Channel. The second order Taylor expansion of (9) leads to following 2nd order term in $\mathcal{K}_{pst}$
[…]
$-(H_{eff,\alpha}\otimes I-I\otimes H_{eff,\alpha}^{\intercal})^{2}$ $=2H_{eff,\alpha}\otimes H_{eff,\alpha}^{\intercal}-H_{eff,\alpha}^{2}\otimes I-I\otimes(H_{eff,\alpha}^{2})^{\intercal}.$ (12)
Since $H_{eff,\alpha}$ is Hermitian this expression has a Lindblad form $\mathcal{L}(A)=A\otimes A^{*}-\frac{1}{2}(A^{\dagger}A)\otimes I-\frac{1}{2}I\otimes(A^{\dagger}A)^{\intercal}$ with a Hermitian dissipator $A=H_{eff,\alpha}$. We conclude that PST nulls the first order of the coherent error and the second order produces a Hermitian Lindbladian. That is
$\mathcal{K}_{PST}$ $\simeq\mathcal{U}_{T}e^{\mathcal{L}_{eff}},$ (13)
$\mathcal{L}_{eff}$ $=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(H_{eff,\alpha}).$ (14)
Since $\mathcal{L}_{eff}=\mathcal{L}_{eff}^{\dagger}$ the resulting noise channel $\mathcal{N}_{PST}=\mathcal{U}_{T}^{\dagger}\mathcal{K}_{PST}\simeq e^{\mathcal{L}_{eff}}$ is also Hermitian up to higher order corrections. We emphasize that unlike RC, $\mathcal{N}_{PST}$ is not a Pauli channel although it is approximately Hermitian. This is easily observed when plotting $\mathcal{N}_{PST}$ in the Pauli basis. The noise map is not diagonal as in RC. This fundamental difference highlights that PST is not a variant of RC or a specific way of implementing RC. It is a different protocol with different analytical features and a different operating regime.
If $H_{coh}$ is a Pauli then $\mathcal{P}_{\alpha}\mathcal{H}_{coh}\mathcal{P}_{\alpha}=\pm\mathcal{H}_{coh}$. However since $\mathcal{L}(A)=\mathcal{L}(-A)$ we get that $\mathcal{L}_{eff}$ simplifies to
[…]
We leave the study of higher order terms to future research.
II.4 PST and Noise
In the context of Pauli twirling, it is evident that the resulting incoherent error exhibits a Pauli error channel structure. However, in the case of PST, this may not hold true, even in the first order. If we repeat the analysis from the previous section but consider a general Lindblad dissipator $\mathcal{L}$ describing a Markovian trace-preserving noise, we obtain
$\mathcal{U}_{T}+\mathcal{U}_{T}\frac{-i}{2^{2n}}\intop\mathcal{U}(t)^{\dagger}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{L}\mathcal{P}_{\alpha}\mathcal{U}(t)dt+O(\mathcal{L}^{2}),$ (16)
[…]
Interestingly, in [35] it was shown that to a leading order, Hermitian noise does not degrade the result of control-free phase estimation. To exploit this feature, the authors applied randomized compiling for transforming the noise into a Pauli channel (which is Hermitian). However this technique can only be applied to Clifford gates. PST can pose an alternative for making the leading order of the noise Hermitian even for short low-noise non-Clifford gates.
III PST features - limitations and opportunities
III.1 PST, Additional Types of Coherent Errors, and Calibration
Coherent errors comprise of either controllable errors arising from miscalibration of controllable degrees of freedom, or uncontrollable errors such as crosstalk that do not depend on the control signals (driving fields). While RC makes no distinction and effectively mitigate both types of coherent errors, PST addresses all coherent error excluding controllable rotation errors (over or under rotation) which play a key role in gate calibration. This feature opens new possibilities.
At first glance, it may seem that this limitation renders PST only a partial solution, as it cannot transform controllable rotation errors into incoherent errors. However, this exact limitation can become an advantage in gate calibration. A proper calibration is always a good practice. When RC is applicable, it will convert the miscalibration error into excessive noise that makes the QEM task more challenging. PST, on the other hand, does not convert controllable rotation errors into incoherent errors. This characteristic allows for the application of PST during the calibration process. By reducing all other coherent errors, and leaving only those associated with the calibration itself, it simplifies the calibration process and improves its accuracy. The high accuracy calibration that PST facilitates, is imperative for circuits with significant depth.
As a concrete scenario we consider a simplified cross-resonance Hamiltonian $H_{d}=a\cos\phi\,\sigma_{z}\otimes\sigma_{x}+a\sin\phi\,\sigma_{z}\otimes\sigma_{y}+\zeta\sigma_{z}\otimes\sigma_{z}$, where $a$ is the amplitude of the drive, $\phi$ is the phase of the drive, and $\zeta$ is the crosstalk coupling between the qubits. See [29] for additional terms. Since $\sigma_{z}\otimes\sigma_{x}$ is the target Hamiltonian, a deviation of the coefficient $a\cos\phi$ from the desired value is considered as a rotation error. Since it is possible to change the sign of this coefficient we consider it as a controllable rotation error. The term $a\sin\phi\,\sigma_{z}\otimes\sigma_{y}$ is a controllable non-rotational error and the last crosstalk term is an uncontrollable non-rotational error. Such scenarios motivate a more formal and accurate classification of coherent errors. Let us assume we have a Hamiltonian of the form
$\mathcal{H}=f(a)\mathcal{H}_{\beta}+\sum_{\gamma\neq\beta}g_{\gamma}(a)\mathcal{H}_{\gamma},$ (17)
where $\mathcal{H}_{\beta}$ is the target Pauli Hamiltonian (multi-qubit Pauli), $\mathcal{H}_{\gamma}$ represents Pauli coherent error terms, and $a$ is the parameter being controlled. Without loss of generality we assume $\mathcal{H}_{\gamma}\neq\mathcal{H}_{\beta}$. If they are equal it is possible to absorb $g(a)$ in the definition of $f(a)$. For brevity, in what follows we consider a single $\mathcal{H}_{\gamma}$ term. Next, we decompose $f$ and $g$ into even and odd contributions, i.e. $f(a)=f_{e}(a)+f_{o}(a)$ where $f_{e}(a)=\frac{1}{2}[f(a)+f(-a)]$ and $f_{o}(a)=\frac{1}{2}[f(a)-f(-a)]$. We refer to the odd terms as controllable since changing the sign of $a$ changes the sign of the terms in the Hamiltonian as needed for PST. Similarly, the even functions represent uncontrollable degrees of freedom from the point of view of PST. Miscalibration of $f(a)$, regardless if the odd or even term is the problem, leads to rotation errors. $g(a)$ stands for non-rotation errors that stems, for example, from crosstalk or parasitic terms in the control signals that generate the target term $\mathcal{H}_{\beta}$. In what follows, it is shown that PST mitigates the terms proportional to $f_{e}(a),g_{o}(a),$ and $g_{e}(a)$, while $f_{o}(a)$ remains unaffected.
While $f_{e}(a),g_{o}(a),$ and $g_{e}(a)$ are small since they represent coherent errors, $f_{o}(a)$ may be potentially large. This is addressed by moving to the interaction picture of $f_{o}(a)\mathcal{H}_{\beta}$. After applying PST using $\mathcal{P_{\alpha}}$ and $a\to sgn(\alpha,\beta)a$, the first order expansion [as in (10)], yields
$\mathcal{K}_{PST}$ $=\mathcal{U}_{T}^{o}+\mathcal{U}_{T}^{o}\frac{-i}{2^{2n}}\intop_{0}^{T}\mathcal{U}^{o}(t)^{\dagger}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\{f_{e}(a)\mathcal{H}_{\beta}+[g_{e}(a)+sgn(\alpha,\beta)g_{o}(a)]\mathcal{H}_{\gamma}\}\mathcal{P}_{\alpha}\mathcal{U}^{o}(t)dt$
$+O(f_{e}^{2},g_{e}^{2},g_{o}^{2},g_{e}g_{o},f_{e}g_{o},f_{e}g_{e}),$
where we have used the parities of $f_{e}$, $g_{e}$ and $g_{o}$ to simplify the expression. $\mathcal{U}^{o}(t)$ is the evolution operator generated by $f_{o}(a)\mathcal{H}_{\beta}$ ($\mathcal{U}_{T}^{o}=\mathcal{U}^{o}(T))$. $\mathcal{U}_{T}^{o}$ contains all potential controlled rotation errors. We can now focus on the effective Hamiltonian:
$\mathcal{H}_{eff}$ $=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\{f_{e}(a)\mathcal{H}_{\beta}+[g_{e}(a)+sgn(\alpha,\beta)g_{o}(a)]\mathcal{H}_{\gamma}\}\mathcal{P}_{\alpha}$ (18)
$=\frac{g_{o}(a)}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}_{\alpha},$ (19)
where we have used the fact that $f_{e}(a)$ and $g_{e}(a)$ do not depend on $\alpha$ and therefore the regular twirling property $\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{H}_{\beta(\gamma)}\mathcal{P}_{\alpha}=0$ nulls these terms. In Appendix II, we show that when $\mathcal{H}_{\gamma}\neq\mathcal{H}_{\beta}$ it holds that $\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}_{\alpha}=0$ and we get $\mathcal{H}_{eff}=0$ so that
$\mathcal{K}_{PST}=\mathcal{U}_{T}^{o}+O(f_{e}^{2},g_{e}^{2},g_{o}^{2},g_{e}g_{o},f_{e}g_{o},f_{e}g_{e}).$ (20)
In the cross-resonance case discussed above $f_{o,zx}(a)=a\cos\phi$, $g_{o,zx}(a)=a\sin\phi$, $f_{e,zx}=0,g_{e,zx}=0$,$g_{o,zz}=0$ and $g_{e,zz}=\zeta$. After PST the effective driving Hamiltonian is
$\mathcal{H}_{d}^{PST}$ $=a\cos\phi\mathcal{H}_{zx}.$ (21)
This leads to a simplification in the calibration procedure: rather than fine-tune $\phi$ to zero and then fine-tune $a$ (e.g. to $\pi/4$), it is possible to coarsely set $\phi\sim 0$ and then calibrate $a$ until $a\cos\phi$ achieves the desired value. Furthermore, PST also suppresses crosstalk to neighboring qubits which improves the accuracy of calibration.
III.2 Alternative Ways for Applying PST in a Circuit
Since PST is versatile and can be applied to any gate or circuit, there are various alternatives for its application within a given circuit. Some options include:
  • Edges PST - Applying PST to the entire circuit, with the twirling gates positioned solely at the circuit’s edges.
  • Gate PST- Applying PST on a per-gate basis.
  • Intra-gate PST - Implementing PST on elements within each gate.
  • Hybrid PST-RC - In Clifford gates, combine Intra-gate PST with RC, or RC with edges PST.
The first three options are illustrated in Fig. 2c. In principle, it appears advantageous to consider all three levels of implementation, or at the very least, the first and third options. Implementing PST in shorter intervals (intra-gate PST) is an effective strategy to prevent the accumulation of coherent errors. Employing PST across the entire circuit is also beneficial, as it can mitigate the accumulation of errors resulting from non-local (non-gate-based) mechanisms.
Intra-Gate PST
Let us elaborate on the intra-gate PST approach. Taking the example of the cross-resonance CNOT gate [29] depicted in Fig. 3a, each non-Clifford element (pulse) can be pseudo twirled independently within the PST framework as shown in Fig. 3b. This contrasts with RC, which cannot be applied to the non-Clifford elements of the gate. Additionally, each one of two segments of the cross-resonance pulse can be divided into multiple shorter pulses, allowing for the application of PST across these smaller segments (Fig. 3c). The prospective advantage here is to mitigate coherent errors before they accumulate. To quantify this claim we show in Appendix III that when the number of slices $m$ is sufficiently large so that neither the drive nor the coherent error changes on the scale of a single slice, the evolution operator is given by:
$\mathcal{K}_{PST}^{(m)}=\mathcal{U}_{T}e^{\frac{T}{m}\intop\mathcal{U}^{\dagger}(t)\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(P_{\alpha}H_{coh}P_{\alpha})\mathcal{U}(t)dt}.$ (22)
While the integral is $m$ independent, the coefficient before it scales like $1/m$ which agrees with the numerical observations in Sec. IV.1. We can quantify the residual noise using the operator norm (see Appendix IV for the explanation why this norm is used)
$\left\|\mathcal{K}_{PST}^{(m)}-\mathcal{U}_{T}\right\|_{op}\simeq\frac{T}{m}\left\|\intop\mathcal{U}^{\dagger}(t)\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(\mathcal{P_{\alpha}}\mathcal{H}_{coh}\mathcal{P_{\alpha}})\mathcal{U}(t)dt\right\|_{op},$ (23)
which clearly scales like $1/m$.
In the context of Clifford gates such as the CNOT, conventional RC can be applied at the gate level, while employing PST within the gate’s internal elements. As PST mitigates the non-rotational coherent errors, RC will have to cope mostly with the residual rotation errors.
Figure 3: (a) A schematic representation of the pulse schedule that generates the IBM echo cross-resonance gate. (b) Intra-gate PST: since PST is not limited to Clifford gates, it can be applied separately to each of the ZX pulses (red lines). (c) Taking it a step further, it is possible to decompose each of the π/4 ZX pulses into $m$ shorter pulses and apply PST to each of these shorter pulses. (d) The slicing effect on a $\pi/2$ ZX pulse. The x-axis represents the number of slices (pulses), while the y-axis indicates the operator norm error from the ideal pulse $\left\|\mathcal{K}_{PST}^{(m)}-\mathcal{U}_{T}\right\|_{op}$. The solid orange line corresponds to a fit using the function $b/m$, where $b$ is a fitting parameter.
III.3 Integration and Compatibility with Quantum Error Mitigation
It is evident that PST, by itself, will only provide a limited solution to errors in quantum devices, as it converts coherent errors into incoherent errors, which still need to be addressed using QEM. Therefore, the impact of PST largely depends on its compatibility with QEM methods. PST was mainly introduced for handling non-Clifford gates. Yet QEM methods that are restricted to Cliffords gates such PEC and PEA may still benefit from intra-gate PST and circuit level PST.
Finally, we point out that much like the ‘Adaptive KIK’ QEM method [24], the feasibility of executing PST hinges on the ability to control the signs of the different terms in the driving Hamiltonian. We have shown in Section II.3 and II.4 that PST turns the non-rotation coherent errors and non Hermitian dissipators into a Hermitian Lindblad dissipator. Hence it is expected that PST will be compatible with the Adaptive KIK method. In Sec. IV.3 and V, we demonstrate numerically and experimentally that PST is not only compatible with Adaptive KIK, but substantially improves the resilience of the method to crosstalk.
IV Numerical Results
IV.1 Intra-Gate PST
As alluded to in Fig. 3c, implementing PST on smaller segments of a pulse allows for further improvement in coherent error suppression. To illustrate this point, we conducted a simulation involving the slicing of a $\pi/2$ cross-resonance pulse. The uncontrolled coherent error is a $\sigma_{z}\otimes\sigma_{z}$ term that accompanies the $\sigma_{z}\otimes\sigma_{x}$ driving Hamiltonian. We have applied PST to each one of the $m$ slices and plotted the error in the resulting evolution operator $\mathcal{K}_{PST}^{(m)}$ with respect to the ideal evolution $\mathcal{U}_{ideal}$ of a $\pi/2$ cross-resonance pulse. Figure 3d shows that the resulting error, which we quantify using $\left\|\mathcal{K}_{PST}^{(m)}-\mathcal{U}_{ideal}\right\|_{op}$, reduces as the number of slices increases. In agreement with (22) we observe that the level of suppression is inversely proportional to the number of slices. The orange curve is a fit of the function $b/m$ to the data, where $b$ is a fit parameter that corresponds to the norm operator of the first Magnus term of the noise [see (23)]. Intuitively, in thin slices the PST mitigates the coherent error before it has a chance to accumulate.
In practice, at some point the duration of the slice becomes comparable to that of the PST Pauli operators, and the small error in these operators sets a performance limit. Nonetheless, our results suggest that before reaching this limiting factor intra-gate PST improves the coherent error suppression.
IV.2 PST as a “Hermitianizer” of the Error Channel
To illustrate the effect of PST on incoherent errors, we consider again a $\pi/2$ echo cross-resonance pulse, but this time with coherent errors replaced by incoherent errors. A $T_{1}$ noise (amplitude damping) is applied to each qubit. To quantify the non-Hermiticity of a noise channel $\mathcal{N}=\mathcal{U}^{\dagger}\mathcal{K}$ we use the measure $G=\left\|\mathcal{N}-\mathcal{N}^{\dagger}\right\|_{op}/\left\|2\mathcal{N}\right\|_{op}$. $G$ equals one for anti-Hermitian operators, and zero for Hermitian operators. Figure 4 displays how the PST suppresses $G$ for different level of initial non-Hermiticity. The initial non-Hermiticity was controlled by varying the strength of the amplitude damping noise. As a result, the Hermitian part of the noise increases as well. We have observed the same behavior when using randomly chosen Lindbladians from an ensemble of random Lindbladians [36].
Figure 4: The effect of PST on incoherent errors (noise). The setup is the same as the cross-resonance pulse setup presented in Fig. 3, but this time with incoherent errors instead coherent errors. The $y$ axis shows the non-Hermiticity suppression factor $G_{no\>PST}/G_{PST}$ where $G=\frac{1}{2}\left\|\mathcal{N}-\mathcal{N}^{\dagger}\right\|_{op}/\left\|\mathcal{N}\right\|_{op}$ measure the non-Hermiticity of the noise channel $\mathcal{N}$.
Figure 5: (a) Integration of pseudo twirling with the Adaptive KIK quantum error mitigation method. Different $x$ values represent the operator norm error of the evolution operator when implementing a 3-qubit transverse Ising time-evolution. The error increase in the $x$ axis is obtained by increasing the coherent error amplitude. The $y$-axis shows the operator norm error after the application of various protocols. The dashed line represents errors without any technique applied. This is a ‘break-even’ line where any result below it represents an improvement. The green curve corresponds to the first-order Adaptive KIK mitigation without PST and the red curve depicts the error reduction when PST is applied without KIK. The combination of KIK and PST (blue curve) showcases a significant error suppression. The inset displays the same plot in a logarithmic scale. For each data point we have used twenty twirlings. (b) Using a hundred twirlings, for initial coherent error of 0.25 (corresponding to the right edge of Fig. (a)) substantially improves the error suppression of KIK+PST as shown by the blue histogram.
IV.3 PST+QEM Integration in the Ising Model
As mentioned earlier, PST alone provides only a partial solution to errors in quantum computers. Ideally, it transforms the noise and coherent errors into a Hermitian noise channel. However, the resulting noise channel can still be too strong to extract meaningful results from a given quantum computer. Therefore, PST must be complemented by a quantum error mitigation protocol. We choose the ‘Adaptive KIK’ QEM method [24] due to its operational simplicity and its compatibility with non-Clifford gates. To exemplify the compatibility of PST with this QEM method, we conducted a three-qubit transverse Ising model simulation with both coherent and incoherent errors. The coherent error Hamiltonian was chosen to be $\{\sigma_{z,i}\}$. Such an error can arise from imprecise determination of the qubits frequencies. Finally, the evolution before the PST is given by:
$\mathcal{K}=e^{-i\,J\sum_{i=1}^{2}\mathcal{H}_{z_{i},z_{i+1}}-i\,g\sum_{i=1}^{3}\mathcal{H}_{x_{i}}-i\,\epsilon\sum_{i=1}^{3}\mathcal{H}_{z_{i}}+\mathcal{L}},$ (24)
where $J$ and $g$ are parameters of the transverse Ising Hamiltonian and $\epsilon$ sets the amplitude of the coherent error. The pulse inverse $\mathcal{K}_{I}$ needed for the KIK protocols is
$\mathcal{K}_{I}=e^{+i\,J\sum_{i=1}^{2}\mathcal{H}_{z_{i},z_{i+1}}+i\,g\sum_{i=1}^{3}\mathcal{H}_{x_{i}}-i\,\epsilon\sum_{i=1}^{3}\mathcal{H}_{z_{i}}+\mathcal{L}}.$ (25)
Importantly, in the pulse inverse $\mathcal{K}_{I}$ the sign of $\epsilon$ remains the same since it represents an uncontrollable coherent error that does not depend on the external drive. The first order KIK mitigated operator is
$\mathcal{K}^{(1)}=\frac{3}{2}\mathcal{K}^{PST}-\frac{1}{2}\mathcal{K}^{PST}\mathcal{K}_{I}^{PST}\mathcal{K}^{PST}.$ (26)
Figure 5(a) illustrates that PST significantly enhances the KIK performance, as evident from the shift from the green curve to the blue curve. To emphasize the insufficiency of PST alone, we have also plotted in red the error when PST is applied without KIK. At the rightmost point the PST+KIK reduces the error by a factor of 65.
In this simulation we randomly sampled the twirling space and used only 20 twirling instances. In the implementation of the evolution $\mathcal{K}$ and $\mathcal{K}\mathcal{K}_{I}\mathcal{K}$ needed for the KIK protocol we used the same twirlings for all $\mathcal{K}$ and $\mathcal{K}_{I}$. Figure 5(b) shows the dependence of the error suppression ratio of the PST combined with 1st order KIK protocol with respect to the original error for $n_{PST}=20$ twirlings (orange) and $n_{PST}=100$ twirlings (blue) . The original noise level is 0.25 (right hand side of Fig. 5(a)). These histograms were generated by repeating the simulation 800 times with different random twirlings. For the simulation parameter we used $J=0.1,g=0.2,\xi=0.00465$. $\epsilon$ was varied from $0.005$ to $0.05$. For the dissipative term $\mathcal{L}$ we use $T_{1}$ noise (amplitude damping) for the different qubits such that $\mathcal{L}=0.5\mathcal{L}_{1}^{T_{1}}+1.7\mathcal{L}_{2}^{T_{1}}+0.3\mathcal{L}_{3}^{T_{1}}$ where the subscript designates the qubit index. The initial state is $\left|000\right\rangle$. The the error was quantified using the operator norm as in Fig. 5.
V Experimental Results
To illustrate the utility of PST to gate-calibration we run a ‘deep calibration’ experiment where we calibrate a sequence of 81 echo cross-resonance (ZX) $\pi/2$ pulses. Such a large number of repetitions amplifies the smallest coherent errors. Yet, a trivial execution of such a long circuit is prone to substantial incoherent errors (noise) that will dramatically effect the calibration curve and lead to an erroneous calibration as demonstrated in [24]. Thus, we use the Adaptive KIK [24] QEM methodology for removing the incoherent errors. We measure the probability to find the system in the initial $\left|00\right\rangle$ state (survival probability, SP) after the 81 repetitions sequence for different stretch factor $\chi$ of the cross-resonance pulse amplitude. $\chi=1$ corresponds to the default amplitude value provided by IBM. In this experiment we use the ibmq_kolkata quantum computer.
By comparing the results of applying KIK with and without PST we observe in Fig. 6 that i) with PST (Fig. 6b) the KIK converges faster (less mitigation orders are needed) compared to the same experiment without PST (Fig. 6a). ii) The value of the calibrated amplitude $\chi(SP=1/2)$ is affected by the PST. iii) Without PST for some $\chi$ values we observe unphysical values. To the best of our knowledge, such non-physical values can occur for two reasons. First, if the mitigation order is too small with respect to the window size parameter in the KIK method (see simulations in the appendix of [24]) nonphysical values can appear. Yet, in this case, it is expected that the unphysical values will disappear in higher mitigation orders. The other reason is pronounced non-Markovian effects. One of the sources of non-Markovian effects in the present setup is crosstalk interaction with adjacent qubits. $n=3$ appears to be more physical than $n=2$, yet, since the two results differ, convergence is not observed and we cannot confirm that unphysical values disappear in higher orders. Moreover, when the mitigation order is too low PST is not expected to help. The fact that PST removes the unphysical values and speeds up convergence suggests that the unphysical values without PST are related to non-Markovian effects. We point out that in other circuits we have suppressed unphysical values without PST by applying dynamical decoupling to neighboring idle qubits.
While we control $\chi$, the actual drive of the ZX term is $\chi a\cos(\phi)$ where $\phi$ is the phase of the drive. In PST calibration it is not needed to know $\phi$ explicitly, $\chi$ is scanned until the survival probability obtain the value of $1/2$ which means that $\chi a\cos(\phi)=\pi/4$ (the value that corresponds to $\pi/2$ cross-resonance pulse).
Figure 6: A ‘deep calibration’ experiment of a cross-resonance gate is conducted on ibmq_kolkata. The initial state is set to $\left|00\right\rangle$, then a sequence of 81 $\pi/2$ echo cross-resonance pulses is executed, aiming at amplifying small coherent errors. The $x$ axis describes the cross-resonance amplitude stretch factor $\chi$. At such depths, the incoherent errors become pronounced, and error mitigation has to be applied. $n=0$ corresponds to no error mitigation, and $n>0$ corresponds to higher-order Adaptive KIK error mitigation methods. The point $\chi(SP=0.5)$, where the graph intersects the value $SP=0.5$ (vertical lines), sets the stretch factor that should be used to recalibrate the amplitude of the cross-resonance pulse. The left figure (a) shows the results without PST, and the right figure (b) shows the results with PST using twenty random twirlings. Not only does PST speeds up the convergence ($n=2$ suffices with PST), but PST also prevents the unphysical results observed without PST (shaded-purple region, probability larger than one). These non-physical results can arise due to non-Markovian crosstalk effects, which PST suppresses. Although the vertical lines indicate that the amplitude should be modified only by $0.2\%$ with respect to the default value set by IBM ($\chi=1$), it is observed that the error in the survival probability is as big as $0.1$ if $\chi=1$ is used instead of $\chi(SP=1/2)\simeq 1.0021$.
One might worry that for $\phi\neq 0$ a residual ZY term will remain and generate a coherent error. However as explained in Sec. III.1, PST eliminates the small $a\sin\phi\,\sigma_{z}\otimes\sigma_{y}$ terms. To experimentally demonstrate this point we carry out a variant of the previous experiment where a $\pi/2$ single-qubit $y$ rotation is added to the target qubit just before the measurement. This makes the survival probability sensitive to $ZY$ rather than to $ZX$. Without PST, it is expected that the outcomes will show some dependence on the amplitude since $\chi a\sin\phi\,\sigma_{z}\otimes\sigma_{y}$ is not exactly zero if $\phi\neq 0$. This behavior is observed in 7a for different mitigation order $n$. However, Fig. 7b demonstrates that after applying PST and KIK we find that the calibration curve is flat as expected. The root of the sum square residuals $\sqrt{SSR}$ in fitting a straight line to the data is presented in the yellow boxes. For $n=3$ the $\sqrt{SSR}$ score with PST is an order of magnitude smaller compared to the no PST case. In this last experiment we used the Taylor coefficients in the KIK method in order to reduce the error bars. Furthermore, for better visibility, in this experiment the sequence contained 41 repetitions of $\pi/2$ echo cross-resonance pulses.
The error of the means $\sigma$ used for plotting the error bars in the PST plot in Fig. 8 and Fig. 9 were calculated using
$\sigma^{2}=\frac{1}{n_{PST}-1}\sum_{j=1}^{n_{PST}}(R_{j}-\frac{1}{n_{PST}}\sum_{k=1}^{n_{PST}}R_{k})^{2},$ (27)
where $n_{PST}$ is the number of PST realizations used, and $R_{j}$ is the mitigated survival probability calculated by averaging over all the shots in a given realizations. At first, it may seem counter intuitive that this simple formula captures correctly the two components of the variance: i) the variance in the measurement of each $R_{k}$ that is determined by the number of shots in each realization, and ii) the variance in the values of $R_{j}$ with respect to its mean ($\frac{1}{n_{PST}}\sum_{k=1}^{n_{PST}}R_{k}$) that is related to the PST twirling of the coherent error. Such a separation to the variance of means and means of variances is possible provided that a good estimation of the means is known [37]. However, our measurements do not provide the exact means but the ‘noisy means’ that includes the additional uncertainty generated by the finite number of shots. We numerically corroborate that the variance of noisy means formula (27) agrees with the variance observed in simulation.
Figure 7: The same experiment as shown in Fig. 6, but this time, instead of measuring the population of the state $\left|00\right\rangle$, we measure the probability of populating the state $\left|0,-x\right\rangle$. This is achieved by applying a $R_{y}(\pi/2)$ rotation before measuring the target qubit. Consequently, the outcome is sensitive to the ZY term instead of the ZX term. Since PST suppresses the residual ZY term and other coherent errors as well, we expect the calibration curve to be flat as the amplitude stretch is varied. As shown in Fig. 7b, the experimental results validate this prediction. In contrast, the $n=3$ curve without PST is not flat. To see this more quantitatively, we calculate the $\sqrt{SSR}$ score (SSR - Sum of Squared Residuals) that indicates the error in a fit to a straight line, and find that it is ten times smaller with PST compared to the same experiment without PST. For clarity, the fitted straight lines are not shown. We used forty random twirling in this experiment.
Finally we point out an important observation regarding the use of virtual Z gates in the pseudo twirling Paulis. Due to their simplicity and zero error rates, virtual Z gates [38] are very popular in quantum computers setups. In short, Z rotations can be pushed further down the circuit by modifying the phases of pulses that generate gates that do not commute with Z. However, despite their advantage, as explained in Appendix V, a virtual Z gate is not suitable for generating the Z gates needed for PST. Essentially, it makes twirlings that should be different become the same, effectively reducing the space of available twirlings. Thus, in our experiment, we implement the $Y_{pulse}$ gate by changing the phase of the $X$ pulse and not by implementing $XZ_{v}$ (where $Z_{v}$ is a virtual Z gate). For $Z$ gates we use $Z_{pulse}=Y_{pulse}X$ or $XY_{pulse}$. See also a similar effect of virtual Z on dynamical decoupling experiments [39].
VI Summary and concluding remarks
The working hypothesis underlying the present paper is that shorter gates facilitate deeper circuits. However, the coherent errors in short multi-qubit gates cannot be addressed with randomized compiling (Pauli twirling) techniques as it hinges on the assumption that the ideal gates are Clifford gates. Our analytical, numerical, and experimental findings indicate that pseudo twirling can bridge this gap and successfully mitigate coherent errors in non-Clifford gates. Pseudo twirling possesses three advantages that separate it from the established randomized compiling approach: i) Applicability to non-Clifford gates: it enables the implementation of significantly shorter gates, resulting in lower noise accumulation compared to Clifford-based implementations. ii) Intra-gate coherent error suppression: pseudo twirling allows for the suppression of coherent errors within gates and sliced pulses, leading to substantial improvement in coherent error mitigation. iii) Edges-PST: by applying pseudo twirling also at the edges of a circuit, it is possible to suppress non-local multi-qubit coherent errors. Moreover, the successful integration of PST with the Adaptive-KIK error mitigation method leads to a substantial performance boost. This boost can be instrumental in achieving quantum advantage in a wide range of fields, including particle physics, condensed matter, quantum chemistry, and more.
Acknowledgements.
The authors acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. We thank IAI/ELTA Systems for the assistance in running the experiments. Raam Uzdin is grateful for support from the Israel Science Foundation (Grant No. 2556/20). The support of the Israel Innovation Authority is greatly appreciated.
Appendix I - Quantum mechanics in Liouville space
Within the conventional Hilbert space framework of Quantum Mechanics, a quantum system of dimension $d$ is described by a density matrix $\rho$ which has the dimensions $d\times d$. In addition, a general completely positive quantum operation can be expressed as
$\rho^{\prime}=\sum_{i}K_{i}\rho K_{i}^{\dagger},$ (28)
where $\rho^{\prime}$ is the transformed density matrix and $\{K_{i}\}$ are Kraus operators that satisfying the relation $\sum_{i}K_{i}^{\dagger}K_{i}=I$ ($I$ being the $d\times d$ identity matrix). In the Hilbert space formalism, observables are described by a Hermitian operators $A$. The expectation value of $A$ for system in the state $\rho$ reads
$\left\langle A\right\rangle=\textrm{Tr}\left(A\rho\right).$ (29)
The Liouville space representation serves as an alternative mathematical formulation that greatly aids in streamlining the expression of quantum states and their transformations. In this model, the density matrix is transformed into a state vector, indicated as $|\rho\rangle$, with a dimensionality of $d^{2}\times 1$, and a generic quantum transformations $\mathcal{O}$ is described by a superoperator matrix with dimensions $d^{2}\times d^{2}$ such that
$|\rho^{\prime}\rangle=\mathcal{\mathcal{O}}|\rho\rangle.$ (30)
Following the methodology of Ref. [40], $|\rho\rangle$ is the column vector whose first $d$ components correspond to the first row of $\rho$, the next $d$ components correspond to the second row of $\rho$, and so forth. More explicitly, the vectorization procedure of a $d\times d$ matrix $B$ reads $|B\rangle=\left(B_{11},B_{12},...,B_{1d},B_{21},B_{22},...,B_{2d},...,B_{d1},B_{d2},...,B_{dd}\right)^{\intercal}$, where $B_{ij}$ is the $ij$ entry of $B$. With this convention, in Liouville space a Kraus map (28) takes the form [40]
$|\rho^{\prime}\rangle=\sum_{i}K_{i}\otimes K_{i}^{\ast}|\rho\rangle,$ (31)
where $K_{i}^{\ast}$ is the element-wise complex conjugate of $K_{i}$. Since unitary maps $\rho^{\prime}=U\rho U^{\dagger}$ are a special case of Kraus map , in Liouville space a unitary evolution is given by
$|\rho^{\prime}\rangle=\mathcal{U}|\rho\rangle=U\otimes U^{\ast}|\rho\rangle.$ (32)
Equation (31) follows from the vectorization rule for a product of three matrices $B$, $C$ and $D$ [40]
$|BCD\rangle=B\otimes D^{\intercal}|C\rangle,$ (33)
where the superscript $t$ denotes transposition (not Hermitian conjugate). Equation (31) follows by setting $B=K_{i}$, $C=\rho$, and $D=K_{i}^{\dagger}$.
Lastly, expectation value (29) of an observable matrix $A_{d\times d}$ can be neatly expressed in Liouville space using the ’bra’ vector $\left\langle A\right|=\left|A\right\rangle^{\dagger}$. That is, as a row vector $\langle A|=\left(A_{11}^{\ast},A_{12}^{\ast},...,A_{1d}^{\ast},...,A_{d1}^{\ast},A_{d2}^{\ast},...,A_{dd}^{\ast}\right)$. Using the Hermiticity of $A$ we obtain that of the expectation value of $A$ can be written as a ’bra-ket’ inner product of $\left|A\right\rangle$ and $\left|\rho\right\rangle$
$\langle A\rangle$ $=\sum_{i,j}A_{ji}\rho_{ij}$
$=\sum_{i,j}A_{ij}^{*}\rho_{ij}$
$=\langle A|\rho\rangle.$ (34)
Appendix II - Derivation of the relation $\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma\neq\beta}\mathcal{P}{}_{\alpha}=0$
[…]
$\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma\neq\beta}\mathcal{P}{}_{\alpha}$ $=\sum_{\alpha\in\beta^{+}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha}+\sum_{\alpha\in\beta^{-}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha},$ (35)
where $\beta^{\pm}$ is the set of $\alpha$’s such that: $\mathcal{P}{}_{\alpha}\mathcal{H}_{\beta}\mathcal{P}{}_{\alpha}=\pm\mathcal{H}_{\beta}$ ($sgn(\alpha,\beta)=\pm 1$). Further simplification leas to
$\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma\neq\beta}\mathcal{P}{}_{\alpha}$ $=\sum_{\alpha\in\beta^{+}}\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha}-\sum_{\alpha\in\beta^{-}}\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha}$
$=2\sum_{\alpha\in\beta^{+}}\mathcal{P}{}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha}.$ (36)
The last equality follows from the twirling property $\sum_{\alpha}\mathcal{P}{}_{\alpha}\mathcal{H}\mathcal{P}{}_{\alpha}=0$ that holds for any $\mathcal{H}$. Next, we define a two-element permutation within $\beta^{+}$. To each element $\mathcal{P}{}_{\alpha_{1}^{+}}$ in $\beta^{+}$ we can assign another Pauli operator $\mathcal{P}{}_{\alpha_{2}^{+}}$ in $\beta^{+}$ using the transformation
$\mathcal{P}{}_{\alpha_{2}^{+}}=\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}},$ (37)
where $\mathcal{P}{}_{\delta}$ is chosen such $\mathcal{P}{}_{\delta}\mathcal{H}_{\beta}\mathcal{P}{}_{\delta}=\mathcal{H}_{\beta}$ and $\mathcal{P}{}_{\delta}\mathcal{H}_{\gamma}\mathcal{P}{}_{\delta}=-\mathcal{H}_{\gamma}$, i.e. $[P_{\delta},P_{\beta}]=0$ and $\{P_{\delta},P_{\gamma}\}=0$. To verify that $\alpha_{2}^{+}\in\beta^{+}$ we check if $\mathcal{P}{}_{\alpha_{2}^{+}}\mathcal{H}_{\beta}\mathcal{P}{}_{\alpha_{2}^{+}}=+\mathcal{H}_{\beta}$
$\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\beta}\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}=\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{P}{}_{\delta}\mathcal{H}_{\beta}\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}=\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\beta}\mathcal{P}{}_{\alpha_{1}^{+}}=+\mathcal{P}{}_{\alpha_{1}^{+}}.$ (38)
which confirms that $\alpha_{2}^{+}\in\beta^{+}$. In the derivation we use the fact that in Liouville space all $\mathcal{P}$ commute with each other. We proceed by studying the outcome of $\mathcal{P}{}_{\alpha_{2}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha_{2}^{+}}$ compared to $\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha_{1}^{+}}$
$\mathcal{P}{}_{\alpha_{2}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha_{2}^{+}}=\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}=\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{P}{}_{\delta}\mathcal{H}_{\gamma}\mathcal{P}{}_{\delta}\mathcal{P}{}_{\alpha_{1}^{+}}=-\mathcal{P}{}_{\alpha_{1}^{+}}\mathcal{H}_{\gamma}\mathcal{P}{}_{\alpha_{1}^{+}}.$ (39)
We conclude that every element $\alpha\in\beta^{+}$ has counterpart element that has opposite sign to the sum (36). As a result, the sum is zero. In the more general case $g_{\gamma}\mathcal{H}_{\gamma}\to\sum_{\gamma}g_{\gamma}\mathcal{H}_{\gamma}$ a different $P_{\delta}$ can be chosen to each element in the sum so the proof above applies automatically to this case as well. Finally, we point out that if it happens that $\{P_{\beta},P_{\gamma}\}=0$ it is possible to choose $P_{\delta}=P_{\beta}$.
Appendix III - derivation of $\mathcal{K}_{pst}^{(m)}$
Under the assumption that the slices are sufficiently short that neither the drive or coherent error change substantially along the slice we can write
$\mathcal{K}_{pst}^{(m)}=e^{-idt\mathcal{H}(t_{1})}e^{+\mathcal{L}_{dt}^{pst}}e^{-idt\mathcal{H}(t_{2})}e^{+\mathcal{L}_{dt}^{pst}}.....e^{-idt\mathcal{H}(t_{m})}e^{+\mathcal{L}_{dt}^{pst}},$ (40)
where $dt=T/m$, $t_{m}=T$, and $\mathcal{L}_{dt}^{pst}=dt^{2}\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(P_{\alpha}H_{coh}P_{\alpha})$ (see Eq. (14)). By inserting the identities $\{U_{k}U_{k}^{\dagger}\}_{k=1}^{m}$ and $U_{m-k}^{\dagger}e^{-idt\mathcal{H}(t_{k+1})}=U_{m-k-1}^{\dagger}$ we obtain:
$\mathcal{K}_{pst}$ $=U_{m}[U_{m}^{\dagger}e^{-idt\mathcal{H}(t_{1})}e^{+\mathcal{L}_{dt}^{pst}}U_{m-1}][U_{m-1}^{\dagger}e^{-idt\mathcal{H}(t_{2})}e^{+\mathcal{L}_{dt}^{pst}}U_{m-2}]$
$\times$ $[U_{m-2}^{\dagger}e^{-idt\mathcal{H}(t_{2})}e^{+\mathcal{L}_{dt}^{pst}}U_{m-3}][U_{m-3}^{\dagger}.....e^{-idt\mathcal{H}(t_{m})}e^{+\mathcal{L}_{dt}^{pst}}$
$=U_{m}[U_{m-1}^{\dagger}e^{+\mathcal{L}_{dt}^{pst}}U_{m-1}][U_{m-2}^{\dagger}e^{+\mathcal{L}_{dt}^{pst}}U_{m-2}][U_{m-3}^{\dagger}e^{+\mathcal{L}_{dt}^{pst}}U_{m-3}]$
$\times$ $U_{m-3}^{\dagger}.....U_{1}U_{1}^{\dagger}e^{-idt\mathcal{H}(t_{m})}e^{+\mathcal{L}_{dt}^{pst}}$
$=U_{m}e^{U_{m-1}^{\dagger}\mathcal{L}_{dt}^{pst}U_{m-1}^{\dagger}}e^{U_{m-2}^{\dagger}\mathcal{L}_{dt}^{pst}U_{m-2}^{\dagger}}e^{U_{m-1}^{\dagger}\mathcal{L}_{dt}^{pst}U_{m-1}^{\dagger}}...e^{+\mathcal{L}_{dt}^{pst}}.$ (41)
By keeping the leading order in $\mathcal{L}_{dt}^{pst}$ we have
$\mathcal{K}_{pst}=\mathcal{U}_{T}e^{\sum_{k}\mathcal{U}_{m-k}^{\dagger}\mathcal{L}_{dt}^{pst}\mathcal{U}_{m-k}}.$ (42)
In regular dissipation $\mathcal{L}_{dt}\propto dt$ so in the limit of large m we obtain the standard first Magnus term. However the present case $\mathcal{L}_{dt}^{PST}\propto dt^{2}=\frac{T}{m}dt$ and therefore additional $T/m$ coefficient appears and we have
$\mathcal{K}_{pst}^{(m)}$ $=\mathcal{U}_{T}e^{\frac{T}{m}\sum_{k=0}^{m-1}\mathcal{U}_{k}^{\dagger}\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(P_{\alpha}H_{coh}P_{\alpha})\mathcal{U}_{k}dt}$
$\to\mathcal{U}_{T}e^{\frac{T}{m}\intop\mathcal{U}^{\dagger}(t)\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(P_{\alpha}H_{coh}P_{\alpha})\mathcal{U}(t)dt}.$ (43)
Appendix IV - using the operator norm for quantifying errors
In our simulations we use the operator norm to describe the error in the evolution operator. In this appendix we show that it is directly related to the error in measuring observables. The operator norm of a matrix $S$ is equal to its largest singular value. Yet, an alternative definition is:
$\left\|S\right\|_{op}=max_{\psi}\frac{|\left\langle\psi\left|S^{\dagger}S\right|\psi\right\rangle|}{\sqrt{\left\langle\psi\left|\psi\right.\right\rangle}}.$ (44)
Applying the Cauchy–Schwarz inequality to the error in some expectation value $\langle A|\rho_{f}\rangle=\langle A|\mathcal{K}|\rho_{0}\rangle$ with respect to the ideal value $\langle A|\mathcal{U}|\rho_{0}\rangle$ we have:
$\left|\left\langle A\left|\mathcal{K}-\mathcal{U}\right|\rho_{0}\right\rangle\right|$ $\leq\sqrt{\left\langle A\left|A\right.\right\rangle}\sqrt{\left\langle\rho_{0}\left|(\mathcal{K}-\mathcal{U})^{\dagger}(\mathcal{K}-\mathcal{U})\right|\rho_{0}\right\rangle}$
$\leq\sqrt{\left\langle A\left|A\right.\right\rangle}\sqrt{\left\langle\rho_{0}\left|\rho_{0}\right.\right\rangle}\left\|\mathcal{K}-\mathcal{U}\right\|_{op},$ (45)
where we used (44) in the last step. Since the constants $\sqrt{\left\langle A\left|A\right.\right\rangle}$ and $\sqrt{\left\langle\rho_{0}\left|\rho_{0}\right.\right\rangle}$ are trivial to evaluate, we conclude that $\left\|\mathcal{K}-\mathcal{U}\right\|_{op}$ sets a clear and easy to calculate bound on the error in any observable $A$ and any initial state $\rho_{0}$. Notably, often the diamond norm is used for comparing different maps. While the diamond norm has some appealing features that can help in proofs, it is difficult to evaluate it when the number of qubits increases. In contrast, the operator norm is trivial to calculate. Furthermore, in the example discussed in Fig. 5, we found that the operator norm and the diamond norm deviate by less than ten percent without any features that distinct one from each other.
Finally, we point out that (45) can be slightly refined when $\mathcal{K}$ is a trace-preserving map. For trace preserving maps $\left\langle I\left|\mathcal{U}\right|\rho_{0}\right\rangle=\left\langle I\left|\mathcal{K}\right|\rho_{0}\right\rangle=1$ and therefore $|\left\langle A\left|\mathcal{K}-\mathcal{U}\right|\rho_{0}\right\rangle|=|\left\langle A+\alpha I\left|\mathcal{K}-\mathcal{U}\right|\rho_{0}\right\rangle|$. This implies that $\sqrt{\left\langle A\left|A\right.\right\rangle}$ can be replaced by $\sqrt{\left\langle A+\alpha I\left|A+\alpha I\right.\right\rangle}$. Minimizing over $\alpha$ we find that
$\left|\left\langle A\left|\mathcal{K}-\mathcal{U}\right|\rho_{0}\right\rangle\right|$ $\leq\sqrt{\left\langle A^{\prime}\left|A^{\prime}\right.\right\rangle}\sqrt{\left\langle\rho_{0}\left|\rho_{0}\right.\right\rangle}\left\|\mathcal{K}-\mathcal{U}\right\|_{op},$ (46)
$A^{\prime}$ $=A-tr(A)I/tr(I).$ (47)
Appendix V - the issue with using virtual gates in PST
The basic idea behind the frame rotation used for virtual Z is as follows. Consider the case where we need to implement an $R_{z}(\theta)$ gate followed by a transformation U. We can write it as:
$Ue^{-\frac{i}{2}\theta\sigma_{z}}\to e^{-\frac{i}{2}\theta\sigma_{z}}e^{+\frac{i}{2}\theta\sigma_{z}}Ue^{-\frac{i}{2}\theta\sigma_{z}}=e^{-\frac{i}{2}\theta\sigma_{z}}[e^{+\frac{i}{2}\theta\sigma_{z}}Ue^{-\frac{i}{2}\theta\sigma_{z}}].$ (48)
For simplicity, we assume that $U=e^{+i\phi P_{\alpha}}$ where $P_{\alpha}$ is a Pauli Matrix. If $[P_{\alpha},\sigma_{z}]=0$, $Ue^{-i\theta\sigma_{z}}=e^{-i\theta\sigma_{z}}U$ and the $e^{-i\theta\sigma_{z}}$ was pushed to the next layer. The idea in virtual gate is to keep pushing the $R_{z}$ gates all the way to the end of the circuit. Since the measurement is in the computational basis, the final $R_{z}$ rotation has no impact.
Now let us consider the more interesting case where $P$ and $\sigma_{z}$ anti-commute:
$e^{+\frac{i}{2}\theta\sigma_{z}}Pe^{-\frac{i}{2}\theta\sigma_{z}}$ $=(\cos\frac{\theta}{2}+i\sin\frac{\theta}{2}\sigma_{z})P(\cos\frac{\theta}{2}-i\sin\frac{\theta}{2}\sigma_{z})$
$=\cos^{2}\frac{\theta}{2}P+i\sin\frac{\theta}{2}[\sigma_{z},P]+\sin^{2}\frac{\theta}{2}\sigma_{z}P\sigma_{z}.$ (49)
Using the anti-commutation:
$(\cos^{2}\frac{\theta}{2}-\sin^{2}\frac{\theta}{2})P+i2\sin\frac{\theta}{2}\cos\frac{\theta}{2}\sigma_{z}P)=\cos\theta P+i\sin\theta\sigma_{z}P.$ (50)
Typically $P$ will be either $\sigma_{x}$ or $\sigma_{y}$. For example, if it is $\sigma_{x}$ we get
$e^{+\frac{i}{2}\theta\sigma_{z}}\sigma_{x}e^{-\frac{i}{2}\theta\sigma_{z}}=\cos\theta\sigma_{x}+i\sin\theta\sigma_{z}\sigma_{x}=\cos\theta\sigma_{x}+\sin\theta\sigma_{y},$ (51)
which means that the physical implementation of $e^{+\frac{i}{2}\theta\sigma_{z}}Pe^{-\frac{i}{2}\theta\sigma_{z}}$ just involves changing the phase of the pulse generating $P_{\alpha}$. For example, an $R_{z}$ before the target qubit of ZX cross-resonance pulse, will manifest in changing the phase of the cross-resonance pulse.
Let us now consider how virtual Z work in PST
$e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\to\mathcal{P}_{IZ}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ}.$ (52)
In pseudo twirling, we use Pauli gates so $\theta=\pi$ in $R_{z}(\theta)$, and therefore the virtualization will amount to changing the sign of the drive without affecting the coherent error
$\mathcal{P}_{IZ_{v}}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ_{v}}=e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ_{v}}\mathcal{P}_{IZ_{v}}=e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}.$ (53)
Thus, a virtual Z twirling is equivalent to the identity twirling and will do nothing in this case. In contrast, when using physical Z gates we get:
$\mathcal{P}_{IZ}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ}=e^{+\mathcal{P}_{IZ}(i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh})\mathcal{P}_{IZ}}=e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{P}_{IZ}\mathcal{H}_{coh}\mathcal{P}_{IZ}},$ (54)
which twirls the coherent error. Similarly, in IY twirling we aim to generate
$e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\to\mathcal{P}_{IY}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IY}.$ (55)
However, if $\mathcal{P}_{IY}$ is generated via virtual Z: $\mathcal{P}_{IY}=\mathcal{P}_{IX}\mathcal{P}_{IZ_{v}}$ we obtain
$\mathcal{P}_{IX}\mathcal{P}_{IZ_{v}}e^{+i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IX}\mathcal{P}_{IZ_{v}}$ $=$
$\mathcal{P}_{IX}e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IZ_{v}}\mathcal{P}_{IX}\mathcal{P}_{IZ_{v}}$ $=$
$\mathcal{P}_{IX}e^{-i\pi/4\mathcal{H}_{ZX}+i\mathcal{H}_{coh}}\mathcal{P}_{IX}.$ (56)
Thus, the IY twirling is indistinguishable from IX twirling. While aiming to sample from the full set of twirlings, the use of virtual Z and also Y using virtual Z, decimates the effective set of accessible twirlings and consequently degrades the ability of PST to mitigate the full scope of uncontrolled coherent errors.